If the principle is 15,300 and the rate is 3.5% and the simple interest is 2,142.00, what is the time, in regards to the missing value?
Answer by addingup(3677)

interest = P.r.t
2142= 15300*0.035*t multiply on left:
2142= 535.50*t Divide both sides by 535.50
2141/535.50= (535.50*t)/535.50 Divide on left, and on right 535.50 multiplying and dividing cancel each other, and you get:
4= t or
t= 4 is the number of periods (if the interest is annual, then 4 years, if the interest is quarterly, then 4 quarters and so on. If the interest is daily then 4 days. You get the idea)